OpenStudy (anonymous):

Find an exact value. sin=(-11pi/12)

OpenStudy (anonymous):

\[ \sin(-x) = -\sin(x) \]

OpenStudy (anonymous):

\[ \sin(x+y) = \sin(x)\cos(y)+\sin(y)\cos(x) \]

OpenStudy (anonymous):

what do you